OpenStudy (anonymous):

If 2x + 3y = 17 and 3x + 4y = 24, then 5x +7y =

OpenStudy (anonymous):

solve for x and y

OpenStudy (anonymous):

x = 4 y = 3 now solve 5x +7y =

OpenStudy (anonymous):

Set up your system of equations so that it is easier to see what is happening. 2x+3y=17 3x+4y=24 You want to manipulate one or the other (or even both) so that a cancellation is possible when you "combine" the equations. -3(2x+3y=17) = -6x-9y=-51 2(3x+4y=24) = 6x+8y= 48 All I did is multiply the top eq by (-3) and multiply the bottom eq by (2). Notice when we "combine" the two resulting equations, the (-6x) and the (6x) give us ZERO. The X variable is gone. Therefore -1y=-3 or y=3. Now you can substitute this value of Y into the first OR the second equation to solve for X. Then you will know both values. You can handle it form here.